# Each Test Image Deserves A Specific Prompt: Continual Test-Time Adaptation for 2D Medical Image Segmentation

### v2

@@ Appendix A Other Implementation Details of Medical Image Segmentation Tasks
GPU device. We conducted all experiments on a single RTX-2080Ti GPU for two medical image segmentation benchmark tasks. Training the source model for the OD/OC segmentation task. We utilized the SGD optimizer with a momentum of 0.99 and a weight decay of 0.0005, and the initial learning rate $lr_{0}$ was set to 0.001 and decayed according to $lr_{t}=lr_{0}\times(1-t/T)^{0.9}$, where $t$ is the current epoch and the maximum epoch $T$ is set to 200. The batch size was set to 8. Empirically, we chose the model trained after the last epoch as the testing model. Training the source model for the polyp segmentation task. We utilized the publicly released code of PraNet [9]https://github.com/DengPingFan/PraNet to train the source model. The Adam algorithm with a learning rate of 0.0001 was adopted as the optimizer and the maximum epoch was set to 20. The batch size was set to 16. Empirically, we chose the model trained after the last epoch as the testing model. Implementation details of different test-time adaptation methods. Since these compared methods (i.e., TENT-continual [44]https://github.com/DequanWang/tent, CoTTA [45]https://github.com/qinenergy/cotta, DLTTA [49]https://github.com/med-air/DLTTA, DUA [27]https://github.com/jmiemirza/DUA, SAR [33]https://github.com/mr-eggplant/SAR, and DomainAdaptor [51]https://github.com/koncle/DomainAdaptor) are all open-source, we adopted their source codes to conduct experiments. We set the batch size to 1 for all experiments following [49].
@@ Appendix B Other Implementation Details of Natural Image Classification Task
GPU device. We conducted all experiments on a single RTX-3090 GPU. Datasets and Evaluation Metrics. PACS [23] dataset is commonly used in domain generalization and test-time adaptation, which comprises 9,991 images and 7 classes that are collected from 4 distinct domains: art, cartoons, photos, and sketches. Training the source model. We trained the source model by empirical risk minimization (ERM) [43] algorithm with ResNet [17] backbone using the SGD optimizer with a learning rate of 5e-5. The batch size was set to 32 and the number of training iteration was set to 5k. We resized all images to $224\times 224$ and used data augmentation during training, including random cropping, random flipping, color jittering, and intensity changing. Implementation details of comparison experiments under the test-time adaptation setup. We compared our VPTTA with four methods (i.e., BN [29], TENT [44], SHOT-IM [25], and TSD [47]) under the test-time adaptation setup following TSD (i.e., given D domains, training on D-1 domains and testing on the left one) and used the publicly released code of TSDhttps://github.com/SakurajimaMaiii/TSD to conduct experiments for all methods. We resized all test images to $224\times 224$ and no data augmentation was used. For all experiments, we set the random seed to 0. To deploy our VPTTA, we utilized the Adam optimizer with a learning rate of 0.01. The hyperparameters $\alpha$ (size of prompt), $S$ (size of memory bank), $K$ (size of support set), and $\tau$ (temperature coefficient in warm-up) are set to 0.02, 64, 4, and 1. The code and weights of pre-trained source models are available at https://github.com/Chen-Ziyang/VPTTA.

### v4

@@ Appendix A Other Implementation Details of Medical Image Segmentation Tasks
GPU device. We conducted all experiments on a single RTX-2080Ti GPU for two medical image segmentation benchmark tasks.
Training the source model for the OD/OC segmentation task. We utilized the SGD optimizer with a momentum of 0.99 and a weight decay of 0.0005, and the initial learning rate $lr_{0}$ was set to 0.001 and decayed according to $lr_{t}=lr_{0}\times(1-t/T)^{0.9}$, where $t$ is the current epoch and the maximum epoch $T$ is set to 200. The batch size was set to 8. Empirically, we chose the model trained after the last epoch as the testing model.
Training the source model for the polyp segmentation task. We utilized the publicly released code of PraNet [9]https://github.com/DengPingFan/PraNet to train the source model. The Adam algorithm with a learning rate of 0.0001 was adopted as the optimizer and the maximum epoch was set to 20. The batch size was set to 16. Empirically, we chose the model trained after the last epoch as the testing model.
Implementation details of different test-time adaptation methods. Since these compared methods (i.e., TENT-continual [44]https://github.com/DequanWang/tent, CoTTA [45]https://github.com/qinenergy/cotta, DLTTA [49]https://github.com/med-air/DLTTA, DUA [27]https://github.com/jmiemirza/DUA, SAR [33]https://github.com/mr-eggplant/SAR, and DomainAdaptor [51]https://github.com/koncle/DomainAdaptor) are all open-source, we adopted their source codes to conduct experiments. We set the batch size to 1 for all experiments following [49].
@@ Appendix B Other Implementation Details of Natural Image Classification Task
GPU device. We conducted all experiments on a single RTX-3090 GPU.
Datasets and Evaluation Metrics. PACS [23] dataset is commonly used in domain generalization and test-time adaptation, which comprises 9,991 images and 7 classes that are collected from 4 distinct domains: art, cartoons, photos, and sketches.
Training the source model. We trained the source model by empirical risk minimization (ERM) [43] algorithm with ResNet [17] backbone using the SGD optimizer with a learning rate of 5e-5. The batch size was set to 32 and the number of training iteration was set to 5k. We resized all images to $224\times 224$ and used data augmentation during training, including random cropping, random flipping, color jittering, and intensity changing.
Implementation details of comparison experiments under the test-time adaptation setup. We compared our VPTTA with four methods (i.e., BN [29], TENT [44], SHOT-IM [25], and TSD [47]) under the test-time adaptation setup following TSD (i.e., given D domains, training on D-1 domains and testing on the left one) and used the publicly released code of TSDhttps://github.com/SakurajimaMaiii/TSD to conduct experiments for all methods. We resized all test images to $224\times 224$ and no data augmentation was used. For all experiments, we set the random seed to 0. To deploy our VPTTA, we utilized the Adam optimizer with a learning rate of 0.01. The hyperparameters $\alpha$ (size of prompt), $S$ (size of memory bank), $K$ (size of support set), and $\tau$ (temperature coefficient in warm-up) are set to 0.02, 64, 4, and 1. The code and weights of pre-trained source models are available at https://github.com/Chen-Ziyang/VPTTA.